OpenStudy (anonymous):

Solve the following system using substitution or elimination: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (king):

y=x+4 x+x+y=46 2x+y=46 2x+x=46 3x=46 x=46/3